# Short-Time Force Response during the Impact of a Droplet with Gas Bubbles

### v1

@@ I Introduction
It may come as no surprise that the above processes have been very well studied experimentally, from the shape the droplets take during flight Villermaux and Bossa (2009), through their spreading Visser et al. (2015), possible splashing, rebouncing, and jetting Xu et al. (2005, 2007); Mou et al. (2021); Zhang et al. (2022), to the force they exert on the surface they fall onto Gordillo et al. (2018); Mitchell et al. (2019). Scientists and engineers have also dedicated their time to study erosion and fracturing effects of droplets impacting solids Bowden and Brunton (1961); Bowden and Field (1964); Field et al. (2012); Kirols et al. (2015), since it is relevant for naval and aviation industry to ensure structural integrity of their vessels cq. aircrafts. Moreover, special attention has been paid to tackle the droplet impact problem using numerical simulations, predominantly with a volume of fluid (VoF) method Visser et al. (2015); Philippi et al. (2016); Zhang et al. (2022); Tretola and Vogiatzaki (2022) or using a boundary integral (BI) approach Davidson (2000); Hicks and Purvis (2010, 2011); Hendrix et al. (2016). This has allowed scientists to access data that is hard to measure experimentally, like the pressure and velocity fields inside and around the droplet.
@@ II.1 Numerical method and lubrication theory
For the bubble inside the droplet, it is assumed that it is filled with a gas that follows an ideal gas Equation Of State (EOS), also called sometimes a polytropic EOS Brennen (2014); Plesset and Prosperetti (1977) such that the pressure inside the bubble $p_{B}$ can be expressed as
@@
where $\mu_{g}$ is the dynamic viscosity of air. After solving 2 for the air radial velocity $u_{r}$, applying the relevant boundary conditions and using continuity, we arrive at an expression that relates the pressure gradient with the vertical position of the bubble $h_{d}(r,t)$, its rate of change $\dot{h}_{d}\equiv\partial h/\partial t$ and the surface velocity of the droplet that is in contact with the bottom air layer $u_{d}$
@@ II.2 Dimensionless parameters
Note that some authors Bouwhuis et al. (2012); Hendrix et al. (2016) relate the capillary number (in the gas phase) to We and St by simply defining $\mathrm{Ca}=\mathrm{We}/\mathrm{St}$. However, this definition is slightly dangerous since one may argue that not $U_{d}$ but the considerably larger $u_{r}$ is the typical velocity in the gas phase. To avoid this problem we decided to characterize our simulations using the Weber and Stokes numbers. The Reynolds number of the droplet, defined as $\mathrm{Re}_{l}=\rho_{l}R_{d}U_{d}/\mu_{l}$, is assumed to be much larger than one, consistent with the use of the inviscid BI method. As for the Reynolds number of the air layer, $\mathrm{Re}_{g}=\rho_{g}h_{d}u_{r}/\mu_{g}$, it is concluded to remain small (order of $10^{-1}$ or smaller) during the entire impact process because of continuity.
@@ III Numerical results
We will now turn to systematically presenting and discussing the results of our numerical simulations. In Fig. 3 we show the variety of systems simulated, in the same order as they will appear in the following subsections. They can be subdivided into two categories, where the first one only involves spherical bubbles inside the droplet (Sec. III.1) and the second in addition to spherical ones also includes toroidal bubbles (Sec. III.2). Since the BI code used is axisymmetrical, any bubble displaced away from the axis of symmetry will automatically become a torus. Figure 3 (a) shows the case where we fix the distance between bubble and droplet’s surface $l$ and vary the bubble’s radius $R$. In (b) we fix $R$ and increase systematically $l$. We then add a vertical bubble array with fixed separation distance between bubbles in (c). After that we continue by adding toroidal bubbles to the system, where the center bubble on the axis of symmetry is always spherical. Since the schematics in Fig. 3 represent cross sections, every torus appears as a pair of circles at the same distance to the left and right from the symmetry axis. In (d) we start adding tori consecutively next to each other with the same distance between them and at the same distance from the droplet’s surface. After filling what we call a row of bubbles, we displace it systematically upwards from the surface, as seen in (e). Finally we simulate several rows of bubbles in the vertical direction as shown in (f).
@@ III.1.1 Fixing $l$, increasing $R$
To determine the origin of this decrease, in Fig. 4 (b) we plot the different pressure profiles $\Delta P(r,t_{f})$ over the solid surface for different bubble sizes, all at the same time $t_{f}=5.52\;\mathrm{\mu s}$, which corresponds to the final time of the NB simulation. First of all we look at the NB case, for which (as in the other cases) the pressure profile presents some distinct features: The pressure in the center (i.e., on the symmetry axis) is minimal, and the profile presents two maxima which are approximately located at the position where the droplet surface has been deformed. This horn-shaped profile (which of course in fact corresponds to the cross section of an axisymmetric structure) is the hallmark of the Wagner profile corresponding to (inviscid) droplet impact without the interference of an intermediate air layer, as discussed below. For now, we observe that in the presence of a bubble both the central value and the maximum value of the pressure are decreased. This can be understood as follows: The bubble can be compressed and deformed, which makes the droplet interface easier to deform as well. As a result the pressure in the center will rise less compared to the NB case, which in the lubrication layer, also affects the maximum pressure. The effect is stronger as the bubble becomes of comparable size to the footprint of the droplet at the touch down point (with a radius of approximately $\sim 100\,\,\mathrm{\mu m}$), whereas for even larger bubbles the result does not significantly change any further. Clearly, the force is smaller for larger bubbles since the integral of $\Delta P$ is smaller for larger $R$. This can be readily seen by neglecting the axisymmetric three-dimensional character and just qualitatively estimating the area under the $\Delta P$ curves.
@@ III.1.2 Fixing $R$, increasing $l$
Figure 6 (a) exhibits the effect of the spherical bubble position on the force. In purple we again plot the NB benchmark case, which presents a higher load for same instances in time and smaller fly time in contrast to the bubbly cases, as previously explained. It is observed that, as the bubble is placed closer to the surface, cushioning is enhanced and the touchdown time increases. Comparing once more the pressure profiles for $t_{f}=5.52\;\mathrm{\mu s}$, it is seen that smaller and slightly wider peaks appear for smaller $l$, Fig. 6 (b). Most significantly, for the smallest distance $l=25$ $\mu$m, the central pressure even becomes negative, which points to compressible deformation of the bubble.
@@ III.2.1 Fixing $l$ and $R$, increasing number of bubbles
In this subsection we investigate the scenario of Fig. 3 (d). The starting case is where a small spherical bubble of $R=25\;\mathrm{\mu m}$ is placed close to the droplet’s surface at $l=10\;\mathrm{\mu m}$, from which we start adding toroidal bubbles to its side for the following simulations. The first torus has a minor radius of $R=25\;\mathrm{\mu m}$ (i.e., the same cross section as the spherical bubble), and the minimum distance between its surface and the droplet is $l=10\;\mathrm{\mu m}$ as well. We apply an initial separation angle of 0.07 rad between bubbles (measured from center to center) in order to place them far enough from one another such that bubbles do not overlap during impact. All other consecutive toroidal bubbles have the same $R$ and $l$ and are again separated by 0.07 rad from each other.
@@ III.2.3 Fixing $l$ and $R$, increasing number of bubble rows
The effect of one full row of bubbles (1S 3T) has been investigated in the last subsection. This final subsection is dedicated to what is depicted in Fig. 3 (f), where more than one, vertically stacked, full rows of bubbles are simulated. The first row is placed at a distance $l=10\;\mathrm{\mu m}$ from the droplet’s surface, and again all bubbles have a radius of $R=25\;\mathrm{\mu m}$ and are separated from each other by 0.07 rad. The second row of bubbles is placed directly above the first one at a distance of $2R+l$ measured from center to center of the bubbles. The third row is placed above the second one at the same separation.
@@ Appendix A Lubrication approximation
In order to obtain an equation for the pressure inside the air layer we start from the radial component of the Navier-Stokes equations in cylindrical coordinates. Using lubrication approximation Oron et al. (1997) we can write
@@
where $\dot{h}_{d}$ is short for $dh_{d}/dt=\partial h_{d}/\partial t+u_{r}\partial h_{d}/\partial r$ at the interface. We can substitute $u_{r}$ from Eq. (9) into (13), perform the integration and multiply by $r$ to find

### v2

@@ I Introduction
It may come as no surprise that the above processes have been very well studied experimentally, from the shape the droplets take during flight Villermaux and Bossa (2009), through their spreading Visser et al. (2015), possible splashing, rebouncing, and jetting Xu et al. (2005, 2007); Mou et al. (2021); Zhang et al. (2022), to the force they exert on the surface they fall onto Gordillo et al. (2018); Mitchell et al. (2019). Also theoretical progress is extensive, starting with the pioneering work of Wagner about the impact of blunt bodies on a liquid surface Wagner (1932) and the pressure profiles these impactors produce onto the liquid, which inspired other seminal articles like the ones from Lesser (1981); Korobkin and Pukhnachov (1988); Scolan and Korobkin (2001); Korobkin and Scolan (2006), or more recently, the work of Philippi et al. (2016), who provided an expression for the pressure profile for the first moments of the impact following a self-similarity argument.
Scientists and engineers have also dedicated their time to study erosion and fracturing effects of droplets impacting solids Bowden and Brunton (1961); Bowden and Field (1964); Field et al. (2012); Kirols et al. (2015), since it is relevant for naval and aviation industry to ensure structural integrity of their vessels and aircrafts. Moreover, special attention has been paid to tackle the droplet impact problem using numerical simulations, predominantly with a volume of fluid (VoF) method Visser et al. (2015); Philippi et al. (2016); Zhang et al. (2022); Tretola and Vogiatzaki (2022) or using a boundary integral (BI) approach Davidson (2000); Hicks and Purvis (2010, 2011, 2013); Hendrix et al. (2016). This has allowed scientists to access data that is hard to measure experimentally, like the pressure and velocity fields inside and around the droplet.
@@ II.1 Numerical method and lubrication theory
For the bubble inside the droplet, it is assumed that it is filled with a gas that follows a polytropic equation of state (EOS) Brennen (2014); Plesset and Prosperetti (1977) such that the pressure inside the bubble $p_{B}$ can be expressed as
@@
where $\mu_{g}$ is the dynamic viscosity of air. After solving 2 for the air radial velocity $u_{r}$, applying the relevant boundary conditions and using continuity, we arrive at an expression that relates the pressure gradient with the vertical position of the droplet $h_{d}(r,t)$, its rate of change $\dot{h}_{d}\equiv\partial h/\partial t$ and the surface velocity of the droplet that is in contact with the bottom air layer $u_{d}$
@@
Note that we use an incompressible formulation of the lubrication layer in (3), whereas one may argue that a compressible formulation would be more appropriate, since we are interested in compressibility effects. One should realize however that the pressure build up in the lubrication layer is dominated by viscous effects regardless of the compressibility of the layer. We have also implemented a compressible formulation of the lubrication layer in our code, but found the code to be less stable and the differences with the incompressible formulation to be negligible for the parameter settings used in this work.
@@ II.2 Dimensionless parameters
Note that some authors Bouwhuis et al. (2012); Hendrix et al. (2016) relate the capillary number (in the gas phase) to We and St by simply defining $\mathrm{Ca}=\mathrm{We}/\mathrm{St}$. However, this definition is slightly dangerous since one may argue that not $U_{d}$ but the considerably larger $u_{r}$ is the typical velocity in the gas phase. To avoid this problem we decided to characterize our simulations using the Weber and Stokes numbers. The Reynolds number of the droplet, defined as $\mathrm{Re}_{l}=\rho_{l}R_{d}U_{d}/\mu_{l}$, is assumed to be much larger than one, consistent with the use of the inviscid BI method. As for the Reynolds number of the air layer, $\mathrm{Re}_{g}=\rho_{g}h_{d}u_{r}/\mu_{g}$, it is concluded to remain small (order of $10^{-1}$ or smaller) during the entire impact process, since, by continuity, $u_{r}h_{d}\sim U_{d}\ell$ with $\ell\sim\sqrt{h_{d}U_{d}}$ we have $\mathrm{Re_{g}}\sim\rho_{g}U_{d}(R_{d}h_{d})^{1/2}/\mu_{g}$, which is a decreasing function of $h_{d}$, i.e., it becomes smaller as $h_{d}$ becomes smaller in time.
Finally, we note that the use of dimensionless quantities is useful if the problem is such that the number of free parameters is reduced considerably by non-dimensionalizing. As we will see in the following section, the parameter space for the simulations is huge, and we could only explore a small part of it. Since, naturally, both representations are equivalent, we chose to continue the discussion using dimensional quantities.
@@ III Numerical results
We will now turn to systematically presenting and discussing the results of our numerical simulations. In Fig. 3 we show the variety of systems simulated, in the same order as they will appear in the following subsections. They can be subdivided into two categories, where the first one only involves spherical bubbles inside the droplet, i.e., bubbles that lie exactly at the axis of symmetry (Sec. III.1) and the second in addition to spherical ones also includes toroidal bubbles (Sec. III.2). Since the BI code used is axisymmetrical, any bubble displaced away from the axis of symmetry will automatically become a torus.
One could argue that the inclusion of toroidal bubbles is dictated by limitations of our numerical model, which is of course true, but arrangements of bubbles entrapped during impact problems are often observed to obey some axial symmetry, even if toroidal bubbles themselves are unstable and would break up quickly by instabilities of the Rayleigh-Plateau or Rayleigh-Taylor type. Such situations have been observed both in experiments (Thoraval et al., 2013) and in the modelling of the impact of partially aerated liquids (Korobkin et al., 2008) or of bodies coated with a thin liquid layer (Korobkin, 2006), where the trapping of air between the bodies was discussed.
Figure 3 (a) shows the case where we fix the distance between bubble and droplet’s surface $l$ and vary the bubble’s radius $R$. In (b) we fix $R$ and increase systematically $l$. We then add a vertical bubble array with fixed separation distance between bubbles in (c). After that we continue by adding toroidal bubbles to the system, where the center bubble on the axis of symmetry is always spherical. Since the schematics in Fig. 3 represent cross sections, every torus appears as a pair of circles at the same distance to the left and right from the symmetry axis. In (d) we start adding tori consecutively next to each other with the same distance between them and at the same distance from the droplet’s surface. After filling what we call a row of bubbles, we displace it systematically upwards from the surface, as seen in (e). Finally we simulate several rows of bubbles in the vertical direction as shown in (f).
For all the aforementioned cases we use a constant number of nodes to describe the surface of each bubble, namely $50$ nodes for spherical bubbles and $100$ for toroidal ones. Since the bubbles can change in volume and deform, the nodes forming them could in theory come very close to each other and eventually overlap, causing the code to stop. In order to avoid this, the nodes are redistributed evenly along natural cubic splines connecting the nodes and describing the bubble surface every second time step.
For the outer droplet surface, we also redistribute the nodes along natural cubic splines every second time step. However in this case, the nodes are not redistributed evenly along the surface but they are allowed to have a minimum and a maximum separation distance, depending on the local droplet curvature. For a droplet of radius 1 (i.e., in dimensionless units), the minimum and maximum node distances are $0.001$ and $0.05$ along the surface, respectively. This was done in order to obtain more accurate results close to the impact zone, where the deformation of the droplet requires a higher node density (smaller node distance), as well as faster simulations since we can simultaneously use less nodes for the upper half of the droplet. Natural cubic splines were used for the redistribution since they smoothly connect the nodes of the free surfaces and they have continuous first and second derivatives.
@@ III.1.1 Fixing $l$, increasing $R$
To determine the origin of this decrease, in Fig. 4 (b) we plot the different pressure profiles $\Delta P(r,t_{f})$ over the solid surface for different bubble sizes, all at the same time $t_{f}=5.52\;\mathrm{\mu s}$, which corresponds to the final time of the NB simulation (see the supplemental material sup for the pressure profiles plotted instead when the minimum separation distance is 350 nm). First of all we look at the NB case, for which (as in the other cases) the pressure profile presents some distinct features: The pressure in the center (i.e., on the symmetry axis) is minimal, and the profile presents two maxima which are approximately located at the position where the droplet surface has been deformed. This horn-shaped profile (which of course in fact corresponds to the cross section of an axisymmetric structure) is the hallmark of the Wagner profile corresponding to (inviscid) droplet impact without the interference of an intermediate air layer, as discussed below. For now, we observe that in the presence of a bubble both the central value and the maximum value of the pressure are decreased. This can be understood as follows: The bubble can be compressed and deformed, which makes the droplet interface easier to deform as well. As a result the pressure in the center will rise less compared to the NB case, which in the lubrication layer, also affects the maximum pressure. The effect is stronger as the bubble becomes of comparable size to the footprint of the droplet at the touch down point (with a radius of approximately $\sim 100\,\,\mathrm{\mu m}$), whereas for even larger bubbles the result does not significantly change any further. Clearly, the force is smaller for larger bubbles since the integral of $\Delta P$ is smaller for larger $R$. This can be readily seen by neglecting the axisymmetric three-dimensional character and just qualitatively estimating the area under the $\Delta P$ curves.
@@ III.1.2 Fixing $R$, increasing $l$
Figure 6 (a) exhibits the effect of the spherical bubble position on the force. In purple we again plot the NB benchmark case, which presents a higher load for same instances in time and smaller fly time in contrast to the bubbly cases, as previously explained. It is observed that, as the bubble is placed closer to the surface, cushioning is enhanced and the touchdown time increases. Comparing once more the pressure profiles for $t_{f}=5.52\;\mathrm{\mu s}$ (see the supplemental material sup for the pressure profiles plotted instead when the minimum separation distance is 350 nm), it is seen that smaller and slightly wider peaks appear for smaller $l$, Fig. 6 (b). Most significantly, for the smallest distance $l=25$ $\mu$m, the central pressure even becomes negative, which points to compressible deformation of the bubble.
@@ III.2.1 Fixing $l$ and $R$, increasing number of bubbles
In this subsection we investigate the scenario of Fig. 3 (d). The starting case is where a small spherical bubble of $R=25\;\mathrm{\mu m}$ is placed close to the droplet’s surface at $l=10\;\mathrm{\mu m}$, from which we start adding toroidal bubbles to its side for the following simulations. The first torus has a minor radius of $R=25\;\mathrm{\mu m}$ (i.e., the same cross section as the spherical bubble), and the minimum distance between its surface and the droplet is $l=10\;\mathrm{\mu m}$ as well. We apply an initial separation angle of 0.07 rad between bubbles (which is the angle between lines joining the centers of the two bubbles to the center of the droplet)) in order to place them far enough from one another such that bubbles do not overlap during impact. All other consecutive toroidal bubbles have the same minor radius $R$, the same minimum distance $l$ to the droplet’s surface and are again separated by an angle of 0.07 rad from each other, i.e., the angle between lines joining the centers of two adjacent bubbles to the center of the droplet is 0.07 rad.
@@
Note that a toroidal bubble with a minor radius $R$ has a substantially larger volume as compared to the center spherical bubble with a radius of the same magnitude. One could choose to run simulations with toroidal bubbles that have the same equivalent volume as the spherical one. However, we decided to take the minor radius of the tori equal to the radius of the spherical bubble in order to have comparable pressures. It should be noted that in our case, the compressibility in both the spherical and the toroidal bubbles is determined by an air layer of approximately the same thickness.
@@ III.2.3 Fixing $l$ and $R$, increasing number of bubble rows
The effect of one full row of bubbles (1S 3T) has been investigated in the last subsection. This final subsection is dedicated to what is depicted in Fig. 3 (f), where more than one, vertically stacked, full rows of bubbles are simulated. The first row is placed at a distance $l=10\;\mathrm{\mu m}$ from the droplet’s surface, and again all bubbles have a radius of $R=25\;\mathrm{\mu m}$ and are separated from each other by 0.07 rad. The second row of bubbles is placed directly above the first row at a distance of $2R+l$ measured from the center of the spherical bubble of the first row to the center of the spherical bubble of the second row. The third row is placed above the second one at the same separation and following the same idea, i.e., the separation distance between the center of the spherical bubble of the second row and the center of the spherical bubble of the third row is $2R+l$.
@@ Appendix A Lubrication approximation
In order to obtain an equation for the pressure inside the air layer we start from the radial component of the incompressible Navier-Stokes equations in cylindrical coordinates. Using lubrication approximation Oron et al. (1997) we can write
@@
where $\dot{h}_{d}$ is short for $\partial h_{d}/\partial t$ at the interface. We can substitute $u_{r}$ from Eq. (9) into (13), perform the integration and multiply by $r$ to find
@@
a) Refer to caption b) Refer to caption c)
Figure 15: Time evolution of the force $F(t)$ that the solid substrate experiences as a droplet of $R_{d}=0.9$ mm (without any bubbles inside) approaches the substrate at a) $U_{d}=0.2$ m/s, b) $U_{d}=2$ m/s and c) $U_{d}=20$ m/s. The cases where the lubrication layer is modelled as incompressible are plotted in blue, and the cases where the lubrication layer is modelled as compressible are plotted in green.
Finally, an analogous compressible formulation of the lubrication layer is equally possible, in which case the problem needs to be supplemented with an equation of state for the gas phase, e.g., expressing adiabatic compression. The numerical problem attached with such an approach, however, is connected to the fact that an additional time derivative of the density appears in the integral in (15), which is depending on the time derivative of the pressure which in turn needs to be obtained from comparing multiple timesteps of the boundary integral part of the simulation. This introduces additional numerical errors that need to be dealt with and substantially increase computation time.
Nevertheless, we have performed such simulations for selected cases and have found differences between the compressible and incompressible approach to be small for the parameter settings used in this experiment, such that we decided to stick with the incompressible approach.
In Fig. 15, we have plotted the time evolution of the force $F(t)$ on the solid when a droplet of $R_{d}=0.9$ mm approaches the substrate at (a) $U_{d}=0.2$ m/s, (b) $U_{d}=2$ m/s and (c) $U_{d}=20$ m/s respectively. The droplet does not contain any bubbles for the aforementioned cases, however we have also run simulations where a bubble is present inside the liquid and found that the difference between incompressible and compressible lubrication layer is minimal for the velocity range explored in this manuscript.